OpenStudy (anonymous):

A circular piece of fabric has a radius of 1.2 feet. The fabric sells for $5.40 per square foot. What is the total cost of the circular piece of fabric? A. $6.48 B. $20.35 C. $24.42 D. $40.69

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

okay thanks so u know what the answer is @plush123

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

40.6944

OpenStudy (igreen):

Find the area.

OpenStudy (anonymous):

im not 100% sure but im 90%sure

OpenStudy (igreen):

\(A = \pi r^2\)

OpenStudy (anonymous):

oh i thought you find the circumference sorry

OpenStudy (igreen):

\(A = (3.14)(1.2^2)\)

OpenStudy (anonymous):

do u agree with that @iGreen

OpenStudy (igreen):

No..

OpenStudy (anonymous):

listen to igreen hes a proffesor

OpenStudy (igreen):

Btw, @plush123 We have to guide him to the answer..and not give it.

OpenStudy (anonymous):

sorry

OpenStudy (igreen):

It's okay

OpenStudy (anonymous):

okay i will @plush123 so what what do i do? @iGreen

OpenStudy (igreen):

Simplify this \(A = (3.14)(1.2^2)\)

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Correct!

OpenStudy (anonymous):

good job

OpenStudy (igreen):

3.14 * 1.2 * 1.2 = 4.5216 4.5216 * 5.4 = 24.41664 24.41664 is approximately $24.42

OpenStudy (anonymous):

THANKS :) @iGreen

OpenStudy (anonymous):

can u help with a one more? @iGreen

OpenStudy (anonymous):

i can

OpenStudy (igreen):

Sure

OpenStudy (anonymous):

okay thank u guys @plush123 @iGreen Question Navigator Choose exactly two answers that are correct. 3. Which figures can be used to find the area of the composite figure? Choose exactly two answers that are correct. A. a triangle with an area of 96 in2 B. a triangle with an area of 192 in2 C. a semicircle with an area of 50 in2 D. a semicircle with an area of 100 in2

OpenStudy (igreen):

One question per post, close this and open a new one.